OpenStudy (ammarah):

Use a sine or cosine ration to find the value of each variable. Round decimals to the nearest tenth.

OpenStudy (ammarah):



OpenStudy (mathstudent55):

In these triangle problems you have to look at the information you are given. First, look at the right angle. The side opposite that angle is the hypotenuse. In this case it's "n". Then look at the angle that is not the right angle that you either have an angle measure or is an unknown you are solving for. Here it's a 50-degree angle. For the 50-degree angle, side "m" is next to it. So side m is adjacent to that angle of 50 degrees. the leg that measures 8 is opposite to the 50-degree angle.

OpenStudy (anonymous):

remembering this might help you out: SohCahToa

OpenStudy (mathstudent55):

To solve for n, keep in mind n is the hypotenuse. For the angle of 50 degrees, the only known side is the leg the measure 8 which is opposite the 50-deg angle. So you have an angle, and opposite side and a hypotenuse. What connects the opposite side to the hypotenuse? It's the sine bec sin A = opp/hyp sin 50 = opp/hyp sin 50 = 8/n n = 8/sin 50 To solve for m, once again, look at the 50-deg angle. The side measuriong 50 is next to it, so it's an adjacent side. You need to connect the adjacent side to the hypotenuse which is n, so you use the cosine function: cos 50 = adj/hyp cos 50 = m/8 m = 8/cos 50

OpenStudy (aravindg):

@mathstudent55 nicely explained :)

OpenStudy (ammarah):

is the answer n=10.5?

OpenStudy (aravindg):

i got 10.44

OpenStudy (aravindg):

n=10.44

OpenStudy (aravindg):

@Chickenisgood delete that u will get banned

OpenStudy (ammarah):

but it says to round to the nearest tenth